What is the name of the sets of numbers of which the square root of 16 belongs, -2.5 belongs, and 7/9 belongs?5 times the sum of 7 and a number. Write an algebraic expression to represent the verbal expression Define a variable,write an equation,and solve the problem: The sum of 3 times a number and 1 is 25. Find the number. The length of a rectangular garden is 7 feet longer than its width. The perimeter of the garden is 38 feet. Find the width and length of the garden. Solve the inequality. -7<9x+2<11,.10d+.05(25-d)>or equal to 1.44. Find the value of K so that the equation below has the solution set {-3}. 3(2x-1)=x(2-k).
Answer by Boreal(15235)

Sum of 3 times a number and 1 is 25
x=number
3x+1=25
3x=24
x=8.
------------------------------
width=x
length=x+7
2(x)+2(x+7) is the perimeter
2x+2x+14=38 feet
4x=24
x=6 feet width
x+7=13 feet length.
-------------------------------
-7<9x+2<11
-9<9x<9
-1 0 should work and -7<2<11
-1 should not work
-7<-7<11; the first part doesn't work
-7<11<11; the second part doesn't work.
----------------------------
.10d+1.25-.05d>=1.44
0.05d>=0.19
d>=3.8
4 should work
.4+0.05(25-3.8)=.4+1.11, and that is greater than 1.44
3 should not work
.3+0.05(22)=.3+1.10 and that is not greater than 1.44.
-------------------------------
5 times the sum of 7 and a number is 5(7+x)
-------------------------------
3(2(-3)-1)=-3(2-k))
3(-7)=-6+3k
-15=3k
k=-5
6x-3=2x+3k
4x=3k+3
4x-3=3k
when k=-5, x= -3